Hi Stuck with this, please help
Find the area under y = cos 2t from t = 0 to t = 0.5.

Thanks Tim

Answer by Alan3354(69443)

Find the area under y = cos 2t from t = 0 to t = 0.5
-----------------
Integral = sin(2t)/2 + C
sin(1)/2 - sin(0)/2 = 0.8414/2 - 0
= 0.4207 sq units